an alloy contains 32% of copper 40% zinc and remaining nickel .find in grams the quantity of each of the content in a sample of 1 kg of alloy
Answer by ankor@dixie-net.com(22740)

an alloy contains 32% of copper 40% zinc and remaining nickel .find in grams the quantity of each of the content in a sample of 1 kg of alloy
:
Find the percent nickel: 100 - 32 - 40 = 28% nickel
1 kg = 1000 grams
.28*1000 = 280 grams of nickel
.32*1000 = 320 grams of copper
.40*1000 = 400 grams of zinc
--------------------------------
check total: 1000 grams which is 1 kg